the probability that a person has received a speeding ticket is 0.35. the probability of a person driving a red car is 0.15. what is the probability of randomly choosing a driver with a speeding ticket who's car is not red.

Answer by ikleyn(52780)

.

Considering the events as independent (which, in the frame of the given problem, is very likelihood), you get

    P(a driver with a speeding ticket who's car is not red) = 0.35*(1-0.15) = 0.35*0.85 = 0.2975.    ANSWER

Solved.

-----------------

There are interesting investigations, supported by insurance companies,
if the car's color and safety driving are dependent or independent events.

You can read about it in the Internet, if you wish.
First Internet article, which came on my request, was this one

https://www.citywidelaw.com/los-angeles-car-accident-attorney/car-color-and-crash-risk/

You may find others.